Please explain how to solve the quadratic equation:
3x^2 + 5x = -2

Answer by stanbon(75887)

Please explain how to solve the quadratic equation:
3x^2 + 5x = -2
-----------------
3x^2 + 5x +2 =0
One solution method is called the AC Method.
---------
Think of two numbers whose product is A*C = 6
and whose sum = B = 5
The numbers are 2 and 3
Replace 5x by 2x + 3x to get:
3x^2+2x+3x+2 = 0
Factor the 1st two and the last two terms separately to get:
x(3x+2)+(3x+2)=0
Factor again:
(3x+2)(x+1)=0
Either 3x+2=0 or x+1=0
Either x = -2/3 or x=-1
===================
Other ways to solve the quadratic are:
graphing
or the quadratic formula
or factoring other than AC Method.
======================
Cheers,
Stan H.